NEW ELECTRONIC SPECTRA OF THE HCCl/DCCl Ã–X̃ VIBRONIC BANDS

CHIA-SHIH LIN, YING-EN CHEN, and \underline{BOR-CHEN CHANG}, Department of Chemistry, National Central University, Chung-Li 32054,Taiwan.

We have successfully acquired new dispersed fluorescence spectra following the excitation of several Ã←X̃ vibronic bands of HCCl and DCCl at visible wavelengths in a discharge free jet expansion using an intensified charge-coupled device (ICCD) detector. The new dispersed fluorescence spectra reveal more details of the X̃^{1}A′ state vibrational structure in these molecules than our previous reports.^{a}^{,}^{b} Complete vibrational parameters including fundamental frequencies, anharmonicities, and coupling constants were determined for the HCCl/DCCl X̃^{1}A′ state. Furthermore, perturbations from the background triplet state (ã^{3}A′′) were clearly observed in the new dispersed fluorescence spectra, and therefore the singlet-triplet energy gap could be determined. Additionally, a couple of new Ã←X̃ vibronic bands were found in the laser excitation spectra of HCCl and DCCl. The rotationally resolved excitation spectra of a few HC^{35}Cl Ã←X̃ vibronic bands and those of HC^{37}Cl were recorded and analyzed. Our progress on the experiments and analyses will be presented.

aC.-W. Chen, T.-C. Tsai, and B.-C. Chang, Chem. Phys. Lett. \underline{347}, 73 (2001).

bC.-L. Lee, M.-L. Liu, and B.-C. Chang, J. Chem. Phys. \underline{117}, 3263 (2002).